# Cardinality-Preserving Structured Sparse Graph Transformers for Molecular Property Prediction

### v1

@@ Abstract
Drug discovery motivates efficient molecular property prediction under limited labeled data. Chemical space is vast (often estimated at $\sim 10^{60}$ drug-like molecules) while only thousands of drugs have been approved [6, 1]. Self-supervised pretraining on large unlabeled corpora has become essential for data-efficient molecular representation learning [10, 3]. We introduce CardinalGraphFormer, a graph transformer that adopts Graphormer-inspired structural biases (SPD and centrality; direct-bond edge bias) within a structured sparse attention regime (SPD $\leq 3$) and augments it with a cardinality-preserving unnormalized aggregation channel over the same support set. Pretraining combines contrastive graph-level alignment with masked attribute reconstruction. Under a fully matched evaluation protocol, CardinalGraphFormer improves mean performance on all 11 tasks and achieves statistically significant gains on 10 of 11 public benchmarks spanning MoleculeNet, OGB, and TDC ADMET tasks over strong reproduced baselines [8, 4, 5].
@@ 1 Introduction
Molecular property prediction serves as a key proxy task in drug discovery, constrained by limited experimental labels and an immense candidate space [6, 1]. Graph transformers with self-supervised pretraining have shown strong performance on standardized benchmarks [9, 10, 3].
We adopt Graphormer-inspired structural biases (SPD + centrality; direct-bond edge bias) but implement structured sparse attention using $K$-hop SPD masks for locality and efficiency. We further augment this with a cardinality-preserving channel to retain dynamic support-size signals complementary to static centrality embeddings.
@@ 2 Related Work
Message passing remains a dominant paradigm for molecular graphs [2], and attention-based GNNs are widely used [7]. Graph transformers such as Graphormer inject structural biases (shortest-path distance, edge features, centrality) into global attention [9]. Sparse variants often restrict attention to local neighborhoods (e.g., $K$-hop masking) to reduce computation in practice. Cardinality-preserving attention adds unnormalized channels to retain degree/cardinality signals [11]. Self-supervised pretraining on graphs includes contrastive methods (GraphCL) and masked modeling (GraphMAE) [10, 3]. CardinalGraphFormer unifies structured sparse attention with aligned cardinality preservation and dual-objective pretraining.
@@ 3.3 Cardinality-Preserving Attention
where $\alpha_{ij}^{(\ell,m)}=\frac{\exp(a_{ij}^{(\ell,m)})}{\sum_{k\in\mathcal{S}(i)}\exp(a_{ik}^{(\ell,m)})}$ and both sums use the head-specific value projections $v_{j}^{(\ell,m)}$ (including self). Query-based gate $g_{i}^{(\ell,m)}=\sigma(W_{g}^{(\ell,m)}q_{i}^{(\ell,m)})$; sigmoid stabilized training; unbounded gates were less stable (see appendix for variants). Concatenation and projection yield $u_{i}^{(\ell)}$, then static centrality injection $\tilde{u}_{i}^{(\ell)}=u_{i}^{(\ell)}+c_{i}$ where $c_{i}=E_{c}(\mathrm{deg\_bin}(i))$. Canonical residual blocks follow (attention + FFN, each with Dropout + Add + LayerNorm). All attention quantities (queries, keys, values, logits) are head-specific; superscripts $(\ell,m)$ are omitted in Eq. 1 for readability.
@@ 3.4 Self-Supervised Objectives
Graph readout: mean-pool final node embeddings; 2-layer MLP projection; NT-Xent ($\tau=0.2$) with in-batch negatives.
@@ 4.4 Results
Table 4: OGB official splits (5 seeds). $\Delta=$ CardinalGraphFormer - SparseGraphormer-K3. significant vs SparseGraphormer-K3.

### v2

@@ Abstract
Drug discovery motivates accurate molecular property prediction when labeled data are limited and candidate spaces are vast. This article presents CardinalGraphFormer, a graph transformer that combines structured sparse attention with graph-specific structural biases and an unnormalized aggregation channel that preserves neighborhood cardinality. The model restricts attention to local shortest-path neighborhoods for efficiency, injects degree-based centrality and direct-bond edge information as biases, and augments normalized attention with a gated sum over the same support set to retain dynamic neighborhood-size signals. Pretraining unifies two self-supervised objectives: contrastive alignment of two augmented graph views and masked reconstruction of node and edge attributes. We evaluate with a matched protocol across multiple public molecular property benchmarks and compare against objective-matched transformer baselines as well as masked-modeling and contrastive baselines. Results show consistent improvements attributable to the cardinality-preserving channel and dual pretraining, alongside ablation evidence and efficiency measurements on large molecules. The work emphasizes reproducibility through shared architectures, data filtering, and split protocols, and discusses limitations tied to sparsity benefits on small graphs.
Keywords: graph transformers; molecular property prediction; self-supervised learning; structured sparsity; cardinality-preserving attention.
@@ 1 Introduction
Molecular property prediction serves as a key proxy task in drug discovery, constrained by limited experimental labels and an immense candidate space [6, 1]. Graph transformers with self-supervised pretraining have shown strong performance on standardized benchmarks [9, 10, 3]. We adopt Graphormer-inspired structural biases (shortest-path distance (SPD) + centrality; direct-bond edge bias) but implement structured sparse attention using $K$-hop SPD masks for locality and efficiency. We further augment this with a cardinality-preserving attention (CPA) channel to retain dynamic support-size signals complementary to static centrality embeddings.
@@ 2 Related Work
Message passing remains a dominant paradigm for molecular graphs [2], and attention-based GNNs are widely used [7]. Graph transformers such as Graphormer inject structural biases (shortest-path distance, edge features, centrality) into global attention [9]. Sparse variants often restrict attention to local neighborhoods (e.g., $K$-hop masking) to reduce computation in practice. Cardinality-preserving attention (CPA) adds unnormalized channels to retain degree/cardinality signals [11]. Self-supervised pretraining on graphs includes contrastive methods (GraphCL) and masked modeling (GraphMAE) [10, 3]. CardinalGraphFormer unifies structured sparse attention with aligned cardinality preservation and dual-objective pretraining.
@@ 3.3 Cardinality-Preserving Attention
where $\alpha_{ij}^{(\ell,m)}=\frac{\exp(a_{ij}^{(\ell,m)})}{\sum_{k\in\mathcal{S}(i)}\exp(a_{ik}^{(\ell,m)})}$ and both sums use the head-specific value projections $v_{j}^{(\ell,m)}$ (including self). Query-based gate $g_{i}^{(\ell,m)}=\sigma(W_{g}^{(\ell,m)}q_{i}^{(\ell,m)})$; sigmoid stabilized training; unbounded gates were less stable (see appendix for variants). Concatenation and projection yield $u_{i}^{(\ell)}$, then static centrality injection $\tilde{u}_{i}^{(\ell)}=u_{i}^{(\ell)}+c_{i}$ where $c_{i}=E_{c}(\mathrm{deg\_bin}(i))$. Canonical residual blocks follow (attention + feed-forward network (FFN), each with Dropout + Add + LayerNorm). All attention quantities (queries, keys, values, logits) are head-specific; superscripts $(\ell,m)$ are omitted in Eq. 1 for readability.
@@ 3.4 Self-Supervised Objectives
Graph readout: mean-pool final node embeddings; 2-layer multilayer perceptron (MLP) projection; NT-Xent ($\tau=0.2$) with in-batch negatives.
@@ 4.4 Results
Table 4: OGB official splits (5 seeds). $\Delta=$ CardinalGraphFormer - SparseGraphormer-K3. denotes paired-bootstrap significance with 95% CI excluding zero vs SparseGraphormer-K3 (Holm-corrected across 11 tasks).